OpenStudy (anonymous):

For what values of t the interval [0,2π] is sin t = √3/2? And explain how to figure out the answer please :)

OpenStudy (amistre64):

draw a unit circle; and the basic trig stuff that they taught you to do lol

OpenStudy (anonymous):

at 60 and 120 degrees. Remember the SOH CAH TOA thing, and memorize the 45 45 90 and 30 60 90 triangle and you're set

OpenStudy (anonymous):

btw, if you don't know it, it goes Sin=Opposite/Hypotenuse Cos=Adjacent/Hypotenuse Tangent=Opposite/Adjacent

OpenStudy (amistre64):

the only way to prove that sin 60 = sqrt(3)/2 is to draw an equilateral triangle, split it in half and pythag it

OpenStudy (amistre64):

then rotate your angle along the unit cirlce thru the quadrants and see which ones are +

OpenStudy (anonymous):

Dur just basic trig. thanks :)

OpenStudy (anonymous):

so, certain angles have exact values as you go from 0 to 2 pi. these are 30= pi/6 , 45= pi/4 and 60= pi/3. every angle similar to it is found by adding or subtracting those values from or to pi or 2 pi. This is called the unit circle, where x = cos of an angle, y= sin of the angle. the sin of pi/3 = sqrt(3)/2 and the cos = (1/2), the cos of pi/6 = sqrt(3) and the sin = 1/2. It is more of a memorization type of thing for thes critical values and how they change. You can find each individual one by drawing a circle with a radius = 1 then using the pythagorean theorem to solve for the trig ratios.